# Gaussian Copula-Based Outage Performance Analysis of Fluid Antenna Systems: Channel Coefficient- or Envelope-Level Correlation Matrix?

### v1

@@ I Introduction
Recently, fluid antenna (FA) has garnered increasing attention in wireless communication communities [1, 2]. Its key idea is to densely deploy ports within a predefined space and dynamically move its RF front-end to the port with the best channel condition for receiving/transmitting signals, thereby increasing spatial diversity. Accordingly, compared to the traditional fixed antenna systems (TAS), FA systems (FAS) enjoy a higher transmission reliability while using much fewer antennas and RF front-ends. However, closely packed ports unavoidably increase spatial correlation among channels of different ports, which poses significant challenges for the performance analysis of FAS.
A rigorous theoretical assessment of FAS performance requires the spatial correlation model to accurately reflect the propagation conditions encountered in practical FAS deployments. To this end, various correlation models have been introduced, including the reference port model in [3], the equally correlated model in [4], the block-diagonal correlation model in [5], and the fully correlated model in [6]. Among these models, the fully correlated model in [6], modeling the channel of each port as a linear combination of channels at other ports, provides the highest accuracy for FAS performance evaluation, however, analytical tractability under this model remains unattainable. To address this issue, copula theory [7] has emerged as an effective approach, which can accurately generate the multivariate distributions of correlated random variables (RVs) by only knowing their marginal distributions, and describe the structure of dependency among multiple RVs through copula function. As a type of elliptical copula, Gaussian copula has attracted considerable attention in the performance analysis of FAS [1, 8], since unlike other types of copulas such as Archimedean copulas, it incorporates a covariance matrix to capture the dependence among multivariate normal RVs. In existing studies, the covariance matrix was approximated in [1] using the channel coefficient correlation matrix from Jake’s model. Accordingly, Gaussian copula links the FAS performance to the number of ports and the size of FA, enabling a more accurate modeling under different port deployment densities compared to other types of copula.
@@ II Gaussian Copula-based Approach for FAS
Next, we compare the modeling accuracy of Gaussian copula when using coefficient- and envelope-level correlation matrices under fully correlated Nakagami-$m$ fading [^2] The PDF and CDF of the empirical Nakagami-$m$ distribution are generated by $f_{|{h^{\left(n\right)}}|}^{{\rm{Nak}}}(r)=\frac{{2{m^{m}}}}{{\Gamma(m){\mu^{m}}}}{r^{2m-1}}{{\rm{e}}^{-\frac{m}{\mu}{r^{2}}}}$ and $F_{|{h^{\left(n\right)}}|}^{{\rm{Nak}}}(r)=\frac{{\gamma\left({m,\frac{m}{\mu}{r^{2}}}\right)}}{{\Gamma(m)}}$, where $\Gamma(\cdot)$ and $\gamma\left({\cdot,\cdot}\right)$ are gamma function and the lower incomplete gamma function, respectively.. Fig. 1 compares the scatterplots of the channel envelopes $\left|{{h^{\left(1\right)}}}\right|$} and $\left|{{h^{\left(2\right)}}}\right|$} following Nakagami-$m$ fading obtained from simulations (i.e., (a), (e), (i)), and under Gaussian copula with the correlation matrix $\bf{R}$} of multivariate normal RVs (i.e., (b), (f), (j)), coefficient-level correlation matrix $\bf{J}$} (i.e., (c), (g), (k)), or the envelope-level correlation matrix $\bf{J}_{h}$} (i.e., (d), (h), (l)). It is worth mentioning that the channel envelope $\left|{{h^{\left(n\right)}}}\right|$ under Gaussian copula can be measured by applying the inverse CDF of Nakagami-$m$ fading to the uniformly distributed RV $u_{n}$, that is $\left|{{h^{\left(n\right)}}}\right|=F_{\left|{{h^{\left(n\right)}}}\right|}^{{\rm{Na}}{{\rm{k}}^{-1}}}\left({{u_{n}}}\right)$. It can be observed that the scatterplots of the channel envelopes under Gaussian copula with the envelope-level correlation matrix and the correlation matrix of multivariate normal RVs are closer to those obtained from simulations, compared to that with the coefficient-level correlation matrix. Specifically, when the port density is high, e.g., $W=0.1$ and $N=2$, the correlation between the channel envelopes generated based on the coefficient-level correlation matrix is stronger than that observed in simulations. This leads to an underestimation of the spatial diversity of FAS and may overestimate the OP. Conversely, when the port density is low, e.g., $W=0.5$ and $N=2$, the channel envelopes generated using the coefficient-level correlation matrix exhibit stronger negative correlation compared to simulations. This results in fewer samples in the lower-left tail of the distribution, which is critical in determining outage events, and may therefore lead to an underestimation of OP. Moreover, although channel envelopes generated using either the correlation matrix of multivariate normal RVs or that of the channel envelopes can better capture the actual sample correlation, little discrepancies still exist particularly at the lower-left tail when the port density is high, potentially introducing performance prediction errors. Such small gap between theory and simulation is expected since the Gaussian copula-based method is an approximation. The modeling accuracy depends not only on the input parameters but also on the chosen copula type. For specific copula type, its accuracy can be improved by selecting appropriate parameters, such as the covariance matrix discussed in this paper. Furthermore, Fig. 2 compares the PDFs and CDFs of $\left|{{h_{{\rm{FAS}}}}}\right|$ under $N=2$, obtained from (1) using coefficient- and envelope-level correlation matrices respectively, as well as from Monte Carlo simulation. In Fig. 2(a), Gaussian copula with envelope-level correlation matrix provides more accurate channel distribution compared to that with coefficient-level one. Moreover, since practical outage requirements are usually below ${\rm{1}}{{\rm{0}}^{-2}}$, i.e., within the CDF value less than ${\rm{1}}{{\rm{0}}^{-2}}$, Fig. 2(b) shows that Gaussian copula with envelope-level correlation matrix can yields a more accurate performance assessment in this regime.
@@ III Construction of Fully Correlated Nakagami-$m$ Fading Channel in FAS
To assess the accuracy of the Gaussian copula-based approach using different correlation matrices, this section presents a method for generating fully correlated Nakagami-$m$ fading channel envelopes to enable accurate performance evaluation via Monte Carlo simulations as a benchmark, and provides the relationship between the coefficient-level and envelope-level correlation matrices.
When generating fully correlated Nakagami-$m$ fading channel envelopes, the following two conditions must be satisfied: 1) The statistical characteristics of the channel envelopes remain unchanged after incorporating the correlation, i.e., the distribution and parameters of the generated envelopes remain unchanged; 2) The correlation between the generated Nakagami-$m$ fading channel envelopes is identical to that between the envelopes derived based on Jake’s model, which will be given in (8).
@@
To satisfy the correlation of channel coefficients, we perform eigenvalue decomposition on the coefficient-level correlation matrix ${\bf{J}}\!=\!{\bf{U\Lambda}}{{\bf{U}}^{\mathcal{H}}}$, where $\mathcal{H}$ stands for conjugate transpose, ${\bf{\Lambda}}\!=\!{\rm{diag}}\left({{\lambda_{1}},\cdots,{\lambda_{N}}}\right)$ is the diagonal eigenvalue matrix with ${\lambda_{1}}\!\!\geq\!\cdots\!\geq\!\!{\lambda_{N}}$, and $\bf{U}$ is the eigenvector matrix. Then, a set of $Nm$ complex Gaussian RVs is generated by
@@
The above relationship between the channel coefficient correlation matrix $\bf{J}$} and the channel gain correlation matrix $\bf{J_{r}}$}}, which is consistent with the relationship specified in [10], verifies the correctness of the proposed channel envelope construction method in (2) and (3). According to [9], the channel envelope correlation matrix can be obtained according to the relationship between the channel gain correlation matrix $\bf{J_{r}}$}} and the channel envelope correlation matrix $\bf{J_{h}}$}}, given by
${\bf{J_{h}}}^{\left({n,\tilde{n}}\right)}=\frac{{F\left({-\frac{1}{2},-\frac{1}{2};m;{\bf{J_{r}}}^{\left({n,\tilde{n}}\right)}}\right)-1}}{{\psi(m)-1}},$ (8)
where $\psi(m)=\Gamma(m)\Gamma(m+1)/{\Gamma^{2}}(m+{1\mathord{\left/{\vphantom{12}}\right.\kern-1.2pt}2})$ and $F(\cdot)$ is the hypergeometric function.
@@
(b) The differences between the correlation matrix of the Nakagami-$m$ channel envelopes generated by different methods.
Figure 3: The distributions and correlation of the Nakagami-$m$ channel envelopes generated by the proposed approach.
To verify the correctness of the proposed fully correlated Nakagami-$m$ channel envelopes construction approach, simulation results in terms of the distribution of the generated channel envelopes and the correlation among them are provided. Fig. 3 (a) shows the comparison of PDFs and CDFs of the generated channel envelopes using the proposed approach and the empirical approach. It can be observed from Fig. 3 (a) that the PDFs and CDFs of the channel envelopes generated by the proposed approach and the empirical approach are perfectly matched. Meanwhile, a quantitative validation based on the root mean square error (RMSE) was conducted, showing that the RMSEs of both the PDFs and CDFs in Fig. 3 (a) under different values of $m$ are all below ${\rm{1}}{{\rm{0}}^{-3}}$. Moreover, Fig. 3 (b) shows the differences between the correlation matrix of the channel envelopes generated by the proposed approach and the channel envelope correlation matrix according (8), and we observe that the differences are very small. These satisfy the above mentioned conditions, proving the feasibility of the proposed fully correlated Nakagami-$m$ channel envelopes construction approach.
@@ IV Numerical Results
(b) The OPs versus the number of ports $N$ for selected values of $\gamma$ when $W=2$.
@@
[…]
Fig. 4 depicts the OPs versus the average transmit SNR $\gamma$ for selected values of $W$ under different methods. Overall, both ‘FAS_Theo_Enve’ and ‘FAS_Theo_Coeff’ can effectively model the outage behavior of FAS, while noticeable discrepancies remain between ‘FAS_Theo_Coeff’ and ‘FAS_Sim’, especially in sparse port deployment scenarios and the low OP regime, compared to the closer match observed with ‘FAS_Theo_Enve’. Specifically, when the size of FA $W$ is large (e.g., $W=3.5$, $N=10$), corresponding to sparse port deployment, ‘FAS_Theo_Enve’ exhibits a clear advantage in modeling accuracy in the low OP regime. When the size of FA $W$ is small (e.g., $W=0.5$, $N=10$), corresponding to dense port deployment, both ‘FAS_Theo_Coeff’ and ‘FAS_Theo_Enve’ yield comparable modeling accuracy.
To further explore the modeling accuracy of Gaussian copula using different correlation matrices in FAS, Fig. 5(a) shows the variation of OPs with respect to the size of FA $W$ for selected values of $N$. It is evident that ‘FAS_Theo_Enve’ offers high accuracy on the whole. In the high-$W$ regime with given $N$ (sparse port deployment), ‘FAS_Theo_Enve’ have an improvement in accuracy compared with ‘FAS_Theo_Coeff’. Moreover, the OPs exhibit fluctuations in the high $W$ regime. This behavior stems from the oscillatory nature of the spatial correlation function ${J_{0}}\left(\cdot\right)$ in the fully correlated model, where the channel of each port is constructed as a weighted combination of others according to the spatial correlation. Furthermore, Fig. 5(b) depicts the OPs versus the number of ports $N$ for selected values of $\gamma$ under different methods. The matching trend further confirms that ‘FAS_Theo_Enve’ can obtain more accurate outage performance than ‘FAS_Theo_Coeff’ in sparse port deployment, i.e., in the low-$N$ regime with given $W$. As the average transmit SNR $\gamma$ increases, ‘FAS_Theo_Enve’ yields more accurate results than ‘FAS_Theo_Coeff’ under denser port deployment, i.e., lager $N$ given a fixed $W$.
@@
Under sparse port deployment, Fig. 6 shows the OPs versus the average transmit SNR $\gamma$ under different methods, for selected values of $W$ and $N$. The discrepancies between ‘FAS_Sim’ and ‘FAS_Theo_Coeff’ as well as ‘FAS_Theo_Coeff’ further validate the above conclusion. Notably, the OPs obtained from ‘FAS_Theo_Coeff’ consistently lie below the accurate OPs obtained from ‘FAS_Sim’, indicating that the Gaussian copula using coefficient-level correlation matrix in existing works underestimates the outage performance of FAS in sparse port deployment scenarios. To explain this phenomenon, Table I compares the resulting channel envelope correlation coefficients (due to the limitation of space, only the channel envelope correlation coefficients between the first port and other ports are presented) obtained from Gaussian copula using the coefficient- and envelope-level correlation matrices. It can be observed from Table I that the correlation coefficients of the generated channel envelope [^3] In Gaussian copula, the resulting channel envelope samples based on different correlation matrices are generated as follows: the correlated uniform samples are first generated using copularnd function in MATLAB with specified Gaussian copula and covariance matrix, and then apply the inverse CDF transformation of Nakagami-$m$ to the generated uniform samples to obtain the corresponding channel envelopes. under Gaussian copula with the envelope-level correlation matrix are much closer to that generated from Monte Carlo simulation. In contrast, using the coefficient-level correlation matrix leads to multiple negative correlation coefficients among the generated channel envelopes. This reduces the probability that multiple channels simultaneously experience deep fading, which in turn underestimates the true outage performance since FAS selects the port with the best channel condition for transmission.
@@
To further explore the impact of Nakagami-$m$ fading parameters on the modeling accuracy of the Gaussian copula when different correlation matrices are used, Fig. 7 plots the OPs versus the average transmit SNR $\gamma$ under different methods, for selected values of $m$ and $\mu$, respectively. On the whole, regardless of the values of $m$ and $\mu$, ‘FAS_Theo_Enve’ consistently provides more accurate OP performance than ‘FAS_Theo_Coeff’ in sparse port deployment and the low OP region. Moreover, as $m$ decreases or $\mu$ increases, ‘FAS_Theo_Enve’ achieves higher accuracy than ‘FAS_Theo_Coeff’ at lower average transmit SNR $\gamma$.

### v3

@@ I Introduction
Recently, fluid antenna (FA) has garnered increasing attention in wireless communication communities [1, 2]. Its key idea is to densely deploy ports within a predefined space and dynamically move its RF front-end to the port with the best channel condition for receiving/transmitting signals, thereby increasing spatial diversity. Accordingly, compared to the traditional fixed antenna systems (TAS), FA systems (FAS) enjoy a higher transmission reliability while using much fewer antennas and RF front-ends. However, closely packed ports unavoidably increase spatial correlation among channels, which poses significant challenges for FAS performance evaluation.
A rigorous theoretical assessment of FAS performance requires the spatial correlation model to accurately reflect the propagation conditions encountered in practical FAS deployments. To this end, various correlation models have been introduced, including the reference port model in [3], the equally correlated model in [4], the block-diagonal correlation model in [5], and the fully correlated model in [6]. Among these models, the fully correlated model in [6] most accurately captures the spatial correlation, where the channel at each port is modeled as a weighted linear combination of the channels at all the other ports, with the weights determined by the eigenvalues and eigenvectors of the correlation matrix. Although the two-stage approximation proposed in [6] results in a simplified OP expression, the analytical framework is not easily extendable to the generalized Nakagami-$m$ fading. Meanwhile, we note that in copula theory [7], the multivariate distribution of correlated random variables (RVs) can be generated from the marginal distributions under arbitrary fading, and that the dependency structure can be modeled using the copula function. As a type of elliptical copula, Gaussian copula has attracted considerable attention in the performance analysis of FAS [1], since unlike other types of copulas such as Archimedean copulas, it incorporates a covariance matrix to capture the dependence among multivariate normal RVs. In existing studies, the covariance matrix was approximated in [1] using the channel coefficient correlation matrix from Jake’s model. Accordingly, Gaussian copula links the FAS performance to the number of ports and size of FA, enabling a more accurate modeling under different port deployment densities compared to other types of copula.
@@ II Gaussian Copula-based Approach for FAS
Next, we compare the modeling accuracy of Gaussian copula when using coefficient- and envelope-level correlation matrices under fully correlated Nakagami-$m$ fading. Fig. 1 compares the scatterplots of the channel envelopes $\left|{{h^{\left(1\right)}}}\right|$} and $\left|{{h^{\left(2\right)}}}\right|$} following Nakagami-$m$ fading obtained from simulations (i.e., (a), (e), (i)), and Gaussian copula with the correlation matrix $\bf{R}$} of multivariate normal RVs [^2] The correlation matrix $\bf{R}$} can be measured as follows. We first generate the Nakagami-$m$ distributed samples and compute their CDF, which follows the uniform distribution. Then, the uniformly distributed samples are mapped as the normally distributed samples through the inverse Gaussian conversion, and then their correlation is calculated. (i.e., (b), (f), (j)), the coefficient-level correlation matrix $\bf{J}$} (i.e., (c), (g), (k)), or the envelope-level correlation matrix $\bf{J}_{h}$} (i.e., (d), (h), (l)). [^3] Under different correlation matrices, multidimensional uniformly distributed samples are generated from a Gaussian copula, and then the envelope samples of $\left|{{h^{\left(n\right)}}}\right|$ can be measured by applying the inverse CDF of Nakagami-$m$ distribution to the uniformly distributed samples. It can be observed that the scatterplots of the channel envelopes under Gaussian copula with the envelope-level correlation matrix and the correlation matrix of multivariate normal RVs are closer to those obtained from simulations, compared to that with the coefficient-level correlation matrix. Specifically, when the port density is high, e.g., $W=0.1$ and $N=2$, the correlation between the channel envelopes generated based on the coefficient-level correlation matrix is stronger than that observed in simulations. This leads to an underestimation of the spatial diversity of FAS and may overestimate the OP. Conversely, when the port density is low, e.g., $W=0.5$ and $N=2$, the channel envelopes generated using the coefficient-level correlation matrix exhibit stronger negative correlation compared to simulations. This results in fewer samples in the lower-left tail of the distribution, which is critical in determining outage events, and may lead to an underestimation of OP. Moreover, although channel envelopes generated using either the correlation matrix of multivariate normal RVs or that of the channel envelopes can better capture the actual sample correlation, little discrepancies still exist particularly at the lower-left tail when the port density is high, potentially introducing performance errors. Such small gap between theory and simulation is expected since the Gaussian copula-based method is an approximation. Furthermore, Fig. 2 compares the PDFs and CDFs of $\left|{{h_{{\rm{FAS}}}}}\right|$ under $N=2$, obtained from (1) using coefficient- and envelope-level correlation matrices respectively, as well as from Monte-Carlo simulation. In Fig. 2(a), Gaussian copula with envelope-level correlation matrix provides more accurate channel distribution compared to that with coefficient-level one. Moreover, since practical outage requirements are usually below ${\rm{1}}{{\rm{0}}^{-2}}$, i.e., within the CDF value less than ${\rm{1}}{{\rm{0}}^{-2}}$, Fig. 2(b) shows that Gaussian copula with envelope-level correlation matrix can yields a more accurate performance assessment in this regime.
@@ III Construction of Fully Correlated Nakagami-$m$ Fading Channel in FAS
This section presents a method for generating fully correlated Nakagami-$m$ fading channel envelopes to enable accurate performance evaluation via Monte-Carlo simulation, which serves as a benchmark to assess the modeling accuracy of Gaussian copula using different correlation matrices.
When generating fully correlated Nakagami-$m$ fading channel envelopes, the following two conditions must be satisfied:
(i) The generated channel envelopes follow Nakagami-$m$ distribution with shape parameter $m$ and scale parameter $\mu$.
(ii) According to [8], the correlation matrix of the generated channel envelopes $\bf{J_{h}}$}} should satisfy
${{\bf{J}}_{\bf{h}}}^{\left({n,\tilde{n}}\right)}=\frac{{F\left({-\frac{1}{2},-\frac{1}{2};m;{{\bf{J}}_{\bf{r}}}^{\left({n,\tilde{n}}\right)}}\right)-1}}{{\psi(m)-1}},\tilde{n}\in{\cal N},n\neq\tilde{n}$, (2)
where $\psi(m)=\Gamma(m)\Gamma(m+1)/{\Gamma^{2}}(m+{1\mathord{\left/{\vphantom{12}}\right.\kern-1.2pt}2})$, $F(\cdot)$ is the hypergeometric function, and ${\bf{J}}_{\bf{r}}$}} denotes the channel gain correlation matrix. Meanwhile, the channel gain correlation matrix ${\bf{J}}_{\bf{r}}$}} should satisfy ${{\bf{J}}_{\bf{r}}}^{\left({n,\tilde{n}}\right)}\!=\!{\left({{{\bf{J}}^{\left({n,\tilde{n}}\right)}}}\right)^{2}}$, as shown in [9].
@@
To satisfy the correlation of channel coefficients, we perform eigenvalue decomposition on the coefficient-level correlation matrix ${\bf{J}}\!=\!{\bf{U\Lambda}}{{\bf{U}}^{\mathcal{H}}}$, where $\mathcal{H}$ stands for conjugate transpose, ${\bf{\Lambda}}\!=\!{\rm{diag}}\left({{\lambda_{1}},\cdots,{\lambda_{N}}}\right)$ is the diagonal eigenvalue matrix with ${\lambda_{1}}\!\!\geq\!\cdots\!\geq\!\!{\lambda_{N}}$, and $\bf{U}$ is the eigenvector matrix. Then, a set of $Nm$ complex Gaussian RVs is generated by [^4] Different from the equally correlated model that employs a common correlation coefficient [8], a correlation matrix $\bf{J}$} is employed under the fully correlated model, which leads to differences in how a set of complex Gaussian RVs is generated for constructing $\left|{{h^{\left(n\right)}}}\right|$ to satisfy the expected correlation.
@@
The above relationship between the channel coefficient correlation matrix $\bf{J}$} and the channel gain correlation matrix $\bf{J_{r}}$}}, which is consistent with the relationship specified in [9], verifies the correctness of the proposed channel envelope construction method in (3) and (4). With obtained $\bf{J_{r}}$}}, the channel envelope correlation matrix ${\bf{J}}_{\bf{h}}$} can be obtained through (2).
@@
(b) The differences between the correlation matrix of the generated envelopes using the proposed approach and the envelope correlation matrix computed by (2).
Figure 3: The distributions and correlation verification of the Nakagami-$m$ channel envelopes generated by the proposed approach.
To verify the correctness of the proposed fully correlated Nakagami-$m$ channel envelopes construction approach, simulation results in terms of the distribution of the generated channel envelopes and the correlation among them are provided. Fig. 3 (a) shows the comparison of PDFs and CDFs of the generated channel envelopes using the proposed approach and the empirical approach. It can be observed from Fig. 3 (a) that the PDFs and CDFs of the channel envelopes generated by the proposed approach and the empirical approach are perfectly matched. Meanwhile, a quantitative validation based on the root mean square error (RMSE) was conducted, showing that the RMSEs of both the PDFs and CDFs in Fig. 3 (a) under different values of $m$ are all below ${\rm{1}}{{\rm{0}}^{-3}}$. Moreover, Fig. 3 (b) shows the differences between the correlation matrix of the channel envelopes generated by the proposed approach and the channel envelope correlation matrix according (2). We can see that the magnitude of differences is below $10^{-2}$}, which is less than $1\%$ with respect to the value of correlation [1,1][-1,1]. These satisfy the above mentioned conditions, proving the feasibility of the proposed fully correlated Nakagami-$m$ channel envelopes construction approach.
@@ IV Numerical Results
(b) The OPs versus the number of ports NN for selected values of $\gamma$ when $W=2$.
@@
[…]
Fig. 4 depicts the OPs versus the average transmit SNR $\gamma$. Overall, both ‘FAS_Theo_Enve’ and ‘FAS_Theo_Coeff’ can effectively model the outage behavior of FAS, while noticeable discrepancies remain between ‘FAS_Theo_Coeff’ and ‘FAS_Sim’, especially in sparse port deployment and the low-OP regime, compared to the closer match observed with ‘FAS_Theo_Enve’. Specifically, when $W$ is large (e.g., $W=3.5$, $N=10$), i.e., sparse port deployment, ‘FAS_Theo_Enve’ exhibits a clear advantage in modeling accuracy in the low-OP regime. When $W$ is small (e.g., $W=0.5$, $N=10$), i.e., dense port deployment, both ‘FAS_Theo_Coeff’ and ‘FAS_Theo_Enve’ yield comparable modeling accuracy.
Fig. 5 (a) shows the variation of OPs with respect to $W$ for selected values of $N$. It is evident that ‘FAS_Theo_Enve’ offers high accuracy on the whole. In the high-$W$ regime with given $N$ (sparse port deployment), ‘FAS_Theo_Enve’ have an improvement in accuracy compared to ‘FAS_Theo_Coeff’. Moreover, the OPs exhibit fluctuations in the high $W$ regime. This behavior stems from the oscillatory nature of ${J_{0}}\left(\cdot\right)$ in the fully correlated model, where the channel at each port is constructed as a weighted combination of that at other ports. Fig. 5 (b) depicts the OPs versus $N$ for selected values of $\gamma$. The matching trend further confirms that ‘FAS_Theo_Enve’ obtains more accurate OP than ‘FAS_Theo_Coeff’ in sparse port deployment, i.e., in the low-$N$ regime with given $W$. As $\gamma$ increases, ‘FAS_Theo_Enve’ yields more accurate results than ‘FAS_Theo_Coeff’ under denser port deployment, i.e., lager $N$ given a fixed $W$.
@@
Under sparse port deployment, Fig. 6 shows the OPs versus $\gamma$ under different methods, for selected values of $W$ and $N$. The discrepancies between ‘FAS_Sim’ and ‘FAS_Theo_Coeff’ as well as ‘FAS_Theo_Coeff’ further validate the above conclusion. Notably, the OPs obtained from ‘FAS_Theo_Coeff’ consistently lie below the accurate OPs obtained from ‘FAS_Sim’, indicating that the Gaussian copula using coefficient-level correlation matrix in existing works underestimates the outage performance of FAS in sparse port deployment. To explain this phenomenon, Table I compares the resulting channel envelope correlation coefficients (due to the limitation of space, only the channel envelope correlation coefficients between the first port and other ports are presented) obtained from Gaussian copula using the coefficient- and envelope-level correlation matrices. It can be observed that the correlation coefficients of the generated channel envelopes based on the envelope-level correlation matrix are much closer to those from Monte-Carlo simulation. However, the coefficient-level correlation matrix produces multiple negative correlation coefficients of the generated channel envelopes. This reduces the probability that multiple channels experience deep fading simultaneously, thereby underestimating the true outage performance of FAS, since the port with the best channel condition is typically selected for transmission. Moreover, low-probability outage events are mainly caused by deep fading of the channels, and thus the OP in the low-outage regime is particularly sensitive to simultaneous deep fading across multiple channels. Therefore, employing the Gaussian copula based on the coefficient-level correlation matrix introduces noticeable modeling errors in the low-outage regime. Besides, as $\gamma$ increases, $\hat{\gamma}$ decreases, and thus the outage events are mainly caused by deep fading of the channels. Therefore, noticeable modeling errors exists for high $\gamma$.
@@
To further explore the impact of Nakagami-$m$ fading parameters on the modeling accuracy of the Gaussian copula when different correlation matrices are used, Fig. 7 plots the OPs versus the average transmit SNR $\gamma$ under different methods, for selected values of $m$ and $\mu$, respectively. On the whole, regardless of the values of $m$ and $\mu$, ‘FAS_Theo_Enve’ consistently provides more accurate OP performance than ‘FAS_Theo_Coeff’ in sparse port deployment and the low OP regime. Moreover, as $m$ decreases or $\mu$ increases, ‘FAS_Theo_Enve’ achieves higher accuracy than ‘FAS_Theo_Coeff’.